OpenStudy (anonymous):

Jake has two similar cylindrical pipes. The radius of the first cylindrical pipe is 5 cm. The circumference of the second cylindrical pipe is 20π cm. The volume of the second cylindrical pipe is how many times greater than the volume of the first cylindrical pipe?

OpenStudy (anonymous):

A) 8 B) 5 C) 3 D) 4

OpenStudy (rajee_sam):

you know the radius of the first one. From the circumference given you can find the radius of the second

OpenStudy (rajee_sam):

Whic is 10 because Circumference = pi x diameter.So diameter is 20 and radius = 10 Now you know the scale factor is 2 so the height of the cylinder will also be twice. Now the formula for volume is\[\pi . r ^{2} . h\] Volume for first Cylinder will be \[\pi . 5 .5. h\] Volume of second Cylinder will be \[\pi . 10 . 10. 2. h\] Now \[\frac{ \pi . 10 . 10. 2. h }{ \pi . 5 . 5. h }\] = 8